x^2 + 2 =2x, then the value of
x^4 -x^3 +x^2 +2 is (explain clearly)
a)0
b)1
c)-1
d)root 2

Answer by stanbon(75887)

If x^2 + 2 = 2x, then the value of
x^4 -x^3 +x^2 +2 is (explain clearly)
------
x^2 = 2x-2
x^4 = (2x-2)^2 = 4x^2 - 8x + 4
x^3 = 2x^2-2x
-----------------------
Ans:: x^4-x^3+x^2+2 = x^4-x^3 + (2x)
Substitute for x^4 and x^3 to get:
= [4x^2 - 8x + 4] - [2x^2-2x] + 2x
= 2x^2 -8x+2x+2x + 4
= 2x^2 - 4x + 4
= 2(x^2 - 2x + 4)
Note: x^2-2x = -2
---
= 2(-2+4)
= 4
Cheers,
Stan H.
--------------
a)0
b)1
c)-1
d)root 2
---------------